The equation of refraction at a spherical surface is


 


μ2 / v – μ1 /u = μ2- μ1 /R


 


taking R=∞, show that this equation leads to equation


 


real depth/ apparent depth = μ21

Grade:12

2 Answers

harini desiraju
13 years ago

when R is infinite, u r talking abt a glass slab.So,u were correct

harini desiraju
13 years ago

when R=infinity,ur actually talking abt a plane interface like in the case of slabs.

(n2/v)-(n1/u)=0

n2/v=n1/u

u/v=n1/n2

as mentioned earlier that a surface with infinite radius will act as a plane interface,'u' corresponds to real depth(object distance) and 'v' corresponds to apparent depth(image distance).Furthur something by infinity is 0.Now check out ur statement.